OpenStudy (yisselluna):

Which of the following would not be the result of a substitution in the following system? 2x + y = 7 y - x = 1 7 - 2x - x = 1 2x + x + 1 = 7 2(y + 1) + y = 7

OpenStudy (masterized):

The answer is B) 2x + x + 1 = 7 Im not 100% sure but I would chose that answer if I was you :) Hope I helped x

OpenStudy (saksh1998):

well all of them can be achieved by substituting x or y in one of the equations except for c so the answer is c A) substitute y=x+1 B) substitute y=x+1 C) substitute x=y-1 but you don't get the required equation fan and medal

jhonyy9 (jhonyy9):

<< fan and medal >> how is this please - ? you wan saying what you wan getting for your help ? but this is opposite to OS rules @mathmale your opinion please about this ? ty.